OpenStudy (anonymous):

what is the probability of rolling a sum of 3 or 5 when rolling two dice

OpenStudy (ankit042):

sum of 3 sum of 5 1,2 3,2 | 4,1 All combination will give =1/6*1/6 now favorable combination = 3*2/36